22f - 10 = 16f - 8
Found 2 solutions by wuwei96815, suresh:
Answer by wuwei96815(245)

In solving equations remember that you can do anything to one side of the equation as long as you do it to the other side also.

22f - 10 = 16f - 8 Add -16f to both sides of the equation.
You get: 22f - 10 - 16f = 16f - 8 - 16f
Simplify: 6f - 10 = - 8 Now, add 10 to each side of the equation.
You get: 6f - 10 + 10 = - 8 + 10
Simplify: 6f = 2 Now, divide both sides by 6.
You get: 6f/6 = 2/6
Simplify: f = 2/6 = 1/3
Does tht make sense to you ?

Answer by suresh(20)

22f - 10 = 16f - 8
Subtract 16f on both sides,
22f - 10 = 16f - 8

-16f = - 16f
---------------
6f - 10 = -8
Add 10 on both sides,
6f - 10 = - 8

+ 10 = +10
--------------
6f = 2
f = 2/6
Hence,the solution is f = 2/6